OpenStudy (anonymous):

ΔXYZ is translated 4 units up and 3 units left to yield ΔX′Y′Z′. What is the distance between any two corresponding points on ΔXYZ and ΔX′Y′Z′ ?

OpenStudy (gorv):



OpenStudy (gorv):

use pythagorse theorm distance=(4^2+3^2)^1/2=(16+9)^1/2=25^1/2=5